4sinx-3cosx = 7 find the number of general solutions of the given expressions.

Grade:11

1 Answers

Arun
6 years ago
Maximum value of any function of form
a sinx -b cosx = \sqrt(a2+ b2 )
Hence
According to given question,
Maximum value is \sqrt(9+16) = 5
Hence this is not possible.
Hence there is no solution.